OpenStudy (anonymous):

Show how to solve: 2 sin^2 x + sin x = 1. Answers are pi/6, 5pi/6, and 3pi/2.

OpenStudy (anonymous):

put it into quadratic form: (2sin -1)(sin+1)=0 Then solve each component for x

OpenStudy (anonymous):

sinx = t 2t^2 +t -1 =0

OpenStudy (anonymous):

t = 1/2, t = -1 sinx = 1/2 sinx =-1